• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Strauch

    Log entry time 08:52:10 on November 05, 2006

    Entry number 188917

    keyword=Raster still not working, taking 12C data

    The raster is still not working. We decided to start data taking on 12C to make use of the 80 uA beam current. We continue to try to contact a technician.